# Assessing Inf luential Observations in Pain Prediction using fMRI Data

### v1

@@ Abstract
Neuroimaging has provided invaluable insight into our understanding of human behavior. Modern analytic approaches combine methodological advances in data acquisition with predictive modeling to map patterns of brain activity onto mental states and behaviors. However, in modeling multivariate patterns of brain activity, the existence of subjects with unusual characteristics could potentially undermine the generalizability of the findings. Attempts to detect and account for these subjects have to date has been sorely lacking in the neuroimgaing domain. In addition, most existing methods for performing inf luential diagnosis, presume a deterministic submodel and are designed for low-dimensional data (i.e., the number of predictors $p$ is smaller than the sample size $n$). However, the stochastic selection of a submodel from high-dimensional data where $p$ exceeds $n$ has become ubiquitous in a wide range of scientific areas including neuroimaging. Therefore, methods for identifying observations that could exert undue inf luence on the choice of a submodel promises to play an important role. To date, discussion of this topic has been limited, falling short in two ways: (i) constrained ability to detect multiple inf luential points, and (ii) applicability only in restrictive settings. After describing the problem, we characterize and formalize the concept of inf luential observations on variable selection. Then, we propose a generalized diagnostic measure, extended from an available metric accommodating different model selectors and multiple inf luential observations, the asymptotic distribution of which is subsequently established for large $p$, thus providing guidelines to ascertain inf luential observations. A high-dimensional clustering procedure is further incorporated into our proposed scheme to detect multiple inf luential points. Simulation is conducted to assess the performances of various diagnostic approaches. The proposed procedure is shown to improve predictive power when analyzing thermal-stimulated pain based on fMRI data.
regression diagnostics,
group deletion,
cluster analysis,
exchangeability,
fMRI,
keywords:
@@
, and
1 Introduction
Neuroimaging is poised to take a substantial leap forward in uncovering the neurophysiological underpinnings of human behavior. This is primarily being driven by the improved analytic and computational capabilities provided by artificial intelligence, machine learning, and associated statistical techniques. This promises the creation of complex brain models, using multivariate patterns of brain activity as input, that will allow us to better understand the functional representations underlying behavior, performance, clinical status and prognosis (Orru et al., 2012; Haynes, 2015). To reach their full potential, it is critical that these models be robust and generalizable across individuals, with strong performance on new, out-of-sample participants. However, the existence of subjects with unusual traits or data could unduly influence the data-dependent formation of the resulting brain models. Therefore, when developing such approaches, it is critical to identify when this occurs to ensure the generalizability of the proposed statistical models.
During the course of a task-based functional magnetic resonance imaging (fMRI) study, a number of subjects receive one or more stimuli while their brain activation is measured at hundreds of time points. The measurement at each time point consists of the subject’s blood oxygenation level dependent (BOLD) response at a large number of spatial locations (voxels), thus giving rise to multivariate time series data. To perform proper statistical analysis at the population-level, it is necessary for voxels to lie in the same location across subjects. The preprocessing of fMRI data therefore includes using nonlinear transformations to transform individual subjects’ anatomical data to a common reference space (e.g., “Montreal Neurologic Institute” (MNI) space). The transformations are thereafter applied to the functional data, allowing for the direct comparison of data across subjects (Lindquist et al., 2008; Ombao et al., 2016). Next, for each subject, a voxel-wise general linear model (GLM) analysis is performed and the estimated regression coefficient corresponding to the task-specific regressor at each voxel is combined into a single subject-specific 3D activation map. This data is used as features in a predictive model. For the model to be generalizable it is important that the data is both comparable across subjects and measuring equivalent phenomena.
As a motivating example, consider the prediction of physical pain using brain activation measures obtained from fMRI data (Lindquist et al., 2017). Pain is associated with large social and economic costs. However, it is primarily assessed via self-report, making it hard to accurately determine. Recently, predictive models have been used to derive patterns of activity across brain regions that provide direct measures of pain intensity (Wager et al., 2013a; Lindquist et al., 2017). However, to be clinically useful these models must exhibit high sensitivity and specificity to pain outcomes. Inter-subject differences in both features and outcomes could be problematic. They could arise for a number of reasons. First, there may exist individual differences in how subjects process pain that influence how different brain regions activate in response to a painful stimulus. Second, though data from all subjects are transformed to a common reference space prior to analysis, there may remain inter-subject variability in functional topology that causes brain locations to be misaligned across subjects (Wang et al., 2022). This could be problematic as brain activation from different locations are often used as features in predictive models, and variation in functional location across subjects will lead to feature misalignment when training models. Third, fMRI data is inherently noisy and prone to artifacts (Lindquist et al., 2008; Ombao et al., 2016). This can lead to the presence of outliers that negatively affect model performance and generalizability.
Together these issues point to the importance of identifying subjects whose data significantly impact the performance of the predictive model. In statistics, the identification of subjects with unusual characteristics or outlying measurements is known as influential diagnosis. This refers to the detection of an isolated subset of the data exhibiting disproportionate influence on various aspects of an estimated model (Belsley et al., 1980). This subset of data is commonly referred to as influential or contaminated observations or data points.
To identify influential observations, it is imperative to first recognize that their detection depends heavily on the particular statistical model being used. In many modern applications, including the fMRI example discussed above, influential point may directly impact model specification as the precise mathematical form of the population is often unspecified beforehand. Indeed, when the number of predictors $p$ substantially surpasses the sample size $n$, high-dimensional variable selection is fundamental to achieve data reduction and attain an interpretable submodel. A non-exhaustive list of such procedures includes the LASSO (Tibshirani, 1996), scaled LASSO (Sun & Zhang, 2012), elastic net (Zou & Hastie, 2005), SCAD (Fan & Li, 2001) and MCP (Zhang, 2010).
Figure 1: Empirical probability of selecting an incorrect submodel based on the contaminated dataset under $(n,p)$= (a) (100,20), and (b) (100,1000). $n_{\text{infl}}$ is the number of influential observations. Details of data generation and perturbation models are given in Section 1 of the Supplementary Materials.
To gain tangible insights into the pitfalls of ignoring observations exerting disproportionate influence on variable selection, a simulated example is given below in Figure 1 with details provided in Section 1 of the Supplementary Materials, schematically illustrating the non-negligible empirical probability of selecting an incorrect submodel based on the full, contaminated dataset. Thus, given the importance and ubiquity of stochastic variable selection, it is necessary and indeed in practitioners’ best interest, to identify observations influencing the choice of a submodel.
Despite its importance, studies on influential diagnostics in model and variable selection is relatively scarce. In the low-dimensional setting, Leger & Altman (1993) pioneered the study by introducing a variant of Cook’s distance (Cook, 1977) incorporating the randomness of model selection for detection purposes. Recently, in the high-dimensional setting, Rajaratnam et al. (2019) proposed and analyzed a diagnostic measure, known as the difference in model selection (DF(LASSO)), to directly quantify the influence of a single point on variable selection via LASSO. In addition, Zhao et al. (2013) studied the so-called high-dimensional influence measure (HIM), capturing the influence of a single observation on the marginal correlation between the response and predictors. The versatility of this measure is then verified on variable selection via simulated and real data studies. This work is extended in Zhao et al. (2015) and Zhao et al. (2019) to detect multiple influential points, leading to the so-called multiple influential point (MIP) detection procedure. Yet, the aforementioned literature falls short on two different fronts:
  1. 1.
    Limitation to detect multiple influential points: the presence of several influential points in a real dataset is common (Hampel et al., 1986). Two major challenges in detecting them are the so-called “masking” and “swamping” effects (Hadi & Simonoff, 1993). Masking is a phenomenon where an influential point is hidden due the presence of other influential cases. Swamping occurs when a non-influential point is falsely labeled as influential. Indeed, under-performance in mitigating them and accurately capturing multiple influential points is seen from the simulated study in Section 4 for both DF(LASSO) and MIP.
  2. 2.
    Restrictive settings: first, as described later in the paper, one claim in developing the DF(LASSO) is not substantiated, jeopardizing its logical legitimacy. Second, the theoretical framework of the DF(LASSO) is only developed for LASSO. Its applicability to other model selectors is yet to be studied. Third, the foregoing works all assume linear regression models. Few studies are conducted for the generalized linear models (GLM).
This manuscript aims at filling the aforementioned gaps and contributes to influential diagnosis on variable selection in five aspects. First, upon characterizing influential points in assessing variable selection, we introduce the notion of sensitivity of model selectors, and draw connections with statistical concepts such as breakdown points and selection consistency. Second, we consolidate the diagnostic measure of Rajaratnam et al. (2019) in Theorem 1 and Corollary 1 by bridging a missing step in its original theoretical development. Third, a generalized version accommodating universal model selectors is then proposed and the corresponding asymptotic distribution is determined in Theorem 2. The main building blocks of our approach in establishing the large sample distribution of our proposed method are the preservation of exchangeability (Commenges, 2003) and characterization of a sequence of arbitrarily dependent events by its exchangeable counterpart (Galambos, 1973). Fourth, upon integrating a high-dimensional clustering procedure into our framework, we propose a novel detection procedure, known as the clustering-based MIP (ClusMIP), that is capable of identifying multiple influential observations. Lastly, we compare our proposal with competing procedures via comprehensive simulated and real data studies. This enables the analysis of strengths and limitations of different procedures, and assessment of sensitivities of model selectors under consideration.
The rest of this manuscript is organized as follows: in Section 2, we characterize and define influential observations on variable selection. Section 3 begins with a survey of the aforementioned HIM and DF(LASSO) approaches for single influential point detection. We then present our generalized version of DF(LASSO) accompanied by the study of its asymptotic properties. We next discuss the MIP and our newly proposed ClusMIP method to detection multiple influential points. Section 4 contains an intensive simulation study, followed by analysis of a real dataset extracted from Lindquist et al. (2017) on the neurologic signature of physical pain in Section 5. Conclusions are given in Section 6.
Notation
@@ 1 Introduction
Let $\text{Y}\in\mathbb{R}$ be the response variable and ${x}_{j}\in\mathbb{R},\,j=1,...,p$, be the predictors. Let $\textbf{Y}=(\text{Y}_{1},...,\text{Y}_{n})^{\top}\in\mathbb{R}^{n}$ be the response vector and $\textbf{X}=(\textbf{X}_{1},...,\textbf{X}_{n})^{\top}\in\mathbb{R}^{n\times p}$, where $\textbf{X}_{i}=(\text{X}_{i1},...,\text{X}_{ip})^{\top}\in\mathbb{R}^{p}$, $i=1,...,n$, be the $n\times p$ design matrix. Then, the $i^{\text{th}}$ observation or point of full dataset $\textbf{Z}\coloneqq(\textbf{Y},\textbf{X})$ is $\text{Z}_{i}\coloneqq(\text{Y}_{i},\textbf{X}_{i}^{\top})^{\top}\in\mathbb{R}^{p+1}$, $i=1,..,n$. Moreover, we abbreviate an arbitrary sequence of random variables or real values $a_{1},...,a_{n}$ by $a_{[n]}$.
@@
2 Assessing Inf luence on Variable Selection
We now characterize and formally define the concept of influential observations on variable selection. It is critical to first recognize that observations influencing the choice of a submodel can only be legitimately understood within a stochastic framework due to the data-dependent nature of model selection. Moreover, due to distinctive model selection criteria, especially different penalty functions of regularized variable selection methods for high-dimensional data, contrasting sets of predictors could be chosen by different model selectors. As such, it is reasonable to conjecture that the extent to which model selectors are perturbed by influential points in terms of selection consistency properties, including model selection consistency (Fan & Lv, 2010) and sign consistency (Zhao & Yu, 2006), varies for different selection procedures. An implication is that influential observations in our context are specific to model selectors, and the same detection procedure as a function of different model selectors may well lead to the detection of distinctive sets of influential points.
In fact, the degree to which a model selector is prone to contamination can be termed as its functional sensitivity to influential observations. This concept can be understood as the percentage of data corruption permitted while fulfilling consistent selection properties. The significance of such sensitivity depends on the dichotomous intents of statistical operations, namely the diagnosis for contaminated data and consistent model selection. For diagnostic purposes, a model selector that is more sensitive to influential points may better facilitate their detection. On the other hand, the later objective of selection consistency underscores the resistance to contamination, which is closely linked to the notion of robustness of an estimator. Under this scenario, functional sensitivity philosophically aligns with the so-called breakdown point (Hampel, 1971; Donoho, 1982) in robust statistics, which is the fraction of data allowed to be perturbed without invalidating the estimation consistency.
From the foregoing discussion, influential points on variable selection can be defined as:
Definition 1.
Having characterized and formally defined influential observations on variable selection, we next discuss methodological developments for detecting them.
3 Methodology
In this section, we present high-dimensional methods for detecting a single and multiple influential observations in Sections 3.1 and 3.2, respectively.
3.1 Single Inf luential Observation Detection
We begin with a concise review of the HIM and DF(LASSO) in Sections 3.1.1 and 3.1.2, respectively. In Section 3.1.3, two theoretical limitations of DF(LASSO) are addressed, followed by the proposal of a generalized version and study of its asymptotic properties.

### v2

@@ Abstract
Neuroimaging data allows researchers to model the relationship between multivariate patterns of brain activity and outcomes related to mental states and behaviors. However, the existence of outlying participants can potentially undermine the generalizability of these models and jeopardize the validity of downstream statistical analysis. To date, the ability to detect and account for participants unduly influencing various model selection approaches have been sorely lacking. Motivated by a task-based functional magnetic resonance imaging (fMRI) study of thermal pain, we propose and establish the asymptotic distribution for a diagnostic measure applicable to a number of different model selectors. A high-dimensional clustering procedure is further combined with this measure to detect multiple inf luential observations. In a series of simulations, our proposed method demonstrates clear advantages over existing methods in terms of improved detection performance, leading to enhanced predictive and variable selection outcomes. Application of our method to data from the thermal pain study illustrates the influence of outlying participants, in particular with regards to differences in activation between low and intense pain conditions. This allows for the selection of an interpretable model with high prediction power after removal of the detected observations. Though inspired by the fMRI-based thermal pain study, our methods are broadly applicable to other high-dimensional data types.
Keywords: regression diagnostics; cluster analysis; exchangeability; fMRI
1 Introduction
Neuroimaging is poised to take a substantial leap forward in uncovering the neurophysiological underpinnings of human behavior. This is primarily being driven by the improved analytic and computational capabilities provided by artificial intelligence, machine learning, and associated statistical techniques. This promises to allow for the creation of complex brain models, using multivariate patterns of brain activity as input, that will allow a better understanding of the functional representations underlying behavior, performance, clinical status, and prognosis [26, 17]. To reach their full potential, it is critical that these models are robust and generalizable across individuals, with strong performance on new, out-of-sample participants. However, the existence of participants with unusual traits or data could unduly influence the formation of the resulting brain models. Therefore, when developing such approaches, it is critical to identify when this occurs to ensure the generalizability of the proposed statistical models.
During the course of a task-based functional magnetic resonance imaging (fMRI) study, participants receive one or more stimuli while their brain activation is measured at hundreds of time points. The measurement at each time point consists of the participant’s blood oxygenation level dependent (BOLD) response at a large number of spatial locations (voxels), thus giving rise to multivariate time series data. To perform a proper statistical analysis at the population level, it is necessary for voxels to lie in the same location across participants. The preprocessing of fMRI data therefore includes using nonlinear transformations to warp individual participants’ anatomical data to a common reference space (e.g., “Montreal Neurologic Institute” (MNI) space). The transformations are then applied to the functional data, allowing for the direct comparison of data across participants [22, 25]. Next, for each participant, a voxel-wise general linear model (GLM) analysis is performed and the estimated regression coefficient corresponding to the task-specific regressor at each voxel is combined into a single subject-specific 3D activation map. This data is often used as input features in a predictive model. For models to be generalizable it is important that the data is both comparable across participants and measures equivalent psychological or physical phenomena. As outlined below, these assumptions may not always hold in practice.
As a motivating example, consider the prediction of physical pain using brain activation maps obtained from fMRI data [21]. Pain is associated with large social and economic costs. However, it is primarily assessed via self-report, making it hard to accurately determine. Recently, predictive models have been used to derive patterns of activity across brain regions that provide direct measures of pain intensity [33, 21]. However, to be clinically useful these models must exhibit high sensitivity and specificity to pain outcomes. Inter-subject differences in both features and outcomes would be problematic and could arise for a number of reasons. First, there may exist individual differences in how participants process pain that influence how different brain regions activate in response to a painful stimulus. For example, not all participants rate a thermal stimulation at a certain temperature as equally painful, and a certain participants may not even find that particular stimulus painful at all. Second, though data from all participants are transformed to a common reference space prior to analysis, there may remain inter-subject variability in functional topology that causes brain locations to be misaligned across participants [34, 35]. This could be problematic as brain activation from different locations are used as features in predictive models, and variation in functional location across participants will lead to feature misalignment when training models. Thus, participants whose functional topology differs substantially from the norm, could unduly influence model performance. Third, fMRI data is inherently noisy and prone to artifacts [22, 25]. Together these three factors can lead to the presence of multiple outliers that negatively affect the formation, performance and generalizability of a brain model. Indeed, models for accessing physical pain could be more susceptible to these issue, as increased pain could potentially cause larger motion related artifacts. This can give rise to drastic changes in BOLD signal, leading to substantial variations in imaging intensity. Thus, the existence of motion outliers in fMRI-based pain studies further impedes the development of a universally valid predictive model, leading to reduced inferential accuracy and reliability.
Together these issues point to the importance of detecting participants whose data significantly impact the performance of the predictive model. In statistics, the detection of participants with unusual characteristics or outlying measurements is known as influential diagnosis. This refers to the detection of an isolated subset of the data exhibiting disproportionate influence on various aspects of an estimated model [3]. This subset of data is commonly referred to as influential (or contaminated) observations or data points.
To identify influential observations, it is imperative to first recognize that their detection depends heavily on the particular statistical model being used. In many modern applications, including the fMRI example discussed above, influential points may directly impact model specification as the precise mathematical form of the population is often unspecified beforehand. Indeed, when the number of predictors $p$ substantially surpasses the sample size $n$, high-dimensional variable selection is a fundamental instrument in exploratory data analysis, designated to achieve data reduction and attain an interpretable submodel. A non-exhaustive list of such procedures includes the LASSO [31], scaled LASSO [29], elastic net [40], SCAD [10] and MCP [36].
To further promote discussion of this topic and gain tangible insights into the pitfalls of ignoring observations exerting disproportionate influence on variable selection, a synthetic numerical example is given in Section 1 of the Supplementary Materials Part I. This illustrates the non-negligible empirical probability of selecting an incorrect submodel based on the full, contaminated dataset. Moreover, this example also explicitly exhibits the adverse impacts of neglecting contaminated observations on the downstream analysis including prediction and model coefficient estimation. Specifically, the mean squared error (MSE) for the predicted responses, the correlation between predicted and true responses, and the MSE for coefficient estimates are all negatively influenced by the presence of influential points. Therefore, given the importance and ubiquity of stochastic variable selection, the possible existence of numerous outliers affecting variable selection especially in fMRI-based pain studies, and the complications caused by ignoring them, it is necessary to identify points influencing the data-dependent choice of a statistical submodel.
Despite its importance, studies on influential diagnostics in model and variable selection is relatively scarce. In addition, to the best of our knowledge there are currently no methods of this type regularly used in neuroimaging. In the low-dimensional setting, [20] pioneered the study by introducing a variant of Cook’s distance [7] incorporating the randomness of model selection for detection purposes. Recently, in the high-dimensional setting, [28] proposed and analyzed a diagnostic measure, known as the difference in model selection (DF(LASSO)), to directly quantify the influence of a single point on variable selection via LASSO. In addition, [37] studied the so-called high-dimensional influence measure (HIM), capturing the influence of a single observation on the marginal correlation between the response and predictors. The versatility of this measure is then verified on variable selection via simulated and real data studies. This work is extended in [38] and [39] to detect multiple influential points, leading to the so-called multiple influential point (MIP) detection procedure.
Yet, the aforementioned literature fall short on several fronts. First, are issues related to detecting multiple influential points. Indeed, multiple outliers are frequently encountered in fMRI-based studies [23]. Their detection is challenging because of “masking” and “swamping” effects [14]. Masking occurs when an influential point is hidden due to the presence of other influential cases, and swamping occurs when a non-influential observation is falsely labeled as influential. Existing methods are based on the leave-one-out scheme, which is not designed to overcome these effects when detecting multiple outliers. This is numerically corroborated in our simulation study (see Figures 1 and 2), as both DF(LASSO) and MIP underperform in capturing multiple influential points. In fact, in simulations resembling the aforementioned thermal pain prediction study, less than $10\%$ of the outliers are recognized by the two existing methods. Due to such inadequate detection, heterogeneity is preserved in the data, leading to high prediction MSE even after outlier removal. Second, a gap exists in the theoretical framework of the DF(LASSO). Moreover, its extension to general penalty functions has not been explored, and the resulting statistical properties have not been carefully established.
This manuscript aims to fill the aforementioned gaps and contribute to influential diagnosis on variable selection in five ways. First, upon characterizing influential points in assessing variable selection, we introduce the notion of sensitivity of model selectors, and draw connections with statistical concepts such as breakdown points and selection consistency. Second, we extend the diagnostic measure of [28] by bridging a missing step in its original theoretical development. Third, a generalized version accommodating universal model selectors is proposed and the corresponding asymptotic distribution determined. The main building blocks in establishing the large sample distribution of our proposed method are the preservation of exchangeability [6] and characterization of a sequence of arbitrarily dependent events by its exchangeable counterpart [13]. Fourth, we propose a novel detection procedure, known as clustering-based MIP (ClusMIP), capable of identifying multiple influential observations. Fifth, we compare our proposal with competing procedures in comprehensive simulation and real data studies. ClusMIP is consistently shown to demonstrate significantly better detection performance than the DF(LASSO). Its application to the scaled LASSO exhibits the highest detection outcome. Consequently, substantial enhancement in prediction and variable selection is observed by removing outliers detected by the ClusMIP. On the other hand, influential points detected by the ClusMIP (applied to MCP) for the fMRI-based pain prediction dataset strengthens the distinction between low and intense pain processing mechanisms. Furthermore, the resulting statistical model is scientifically interpretable and displays the best prediction performance after outlier removal.
The manuscript is organized as follows: Section 2 describes the motivating fMRI dataset [21] used to create predictive model for physical pain. Section 3 characterizes and defines influential points on variable selection. Section 4 begins with a survey of the existing HIM and DF(LASSO) approaches for single influential point detection. We then present our generalized version called the generalized difference in model selection (GDF) measure accompanied by the study of its asymptotic properties. We next discuss the MIP and propose our clustering-based detection scheme known as the ClusMIP to identify multiple influential observations. Section 5 contains an intensive simulation study, followed by analysis of the fMRI dataset in Section 6. Conclusions and further discussions are given in Section 7.
2 An fMRI Study of Thermal Pain
@@
The structural T1-weighted images were co-registered to the mean functional image for each participant using the iterative mutual information-based algorithm implemented in the statistical parametric mapping (SPM) software package [2], and were then normalized to the Montreal Neurological Institute (MNI) space. In each functional dataset, initial volumes were removed to allow for stabilization of image intensity. Prior to processing functional images, we removed volumes with signal values that were outliers within the time series. This was done by first computing the mean and the standard deviation of intensity values across each slice, for each image, and then computing the Mahalanobis distances for the matrix of slice-wise mean and standard deviation values. Values with a significant $\chi^{2}$-value (corrected for multiple comparisons based on the more stringent of either false discovery rate or Bonferroni methods) were considered outliers. Next, functional images were slice-time corrected and motion-corrected using SPM. Functional images were warped to SPM’s normative atlas (warping parameters estimated from co-registered, high-resolution structural images) and interpolated to 2 $\times$ 2 $\times$ 2 $\text{mm}^{3}$ voxels.
Previous work on data using a similar experimental paradigm have shown significant individual differences in how participants process pain that influence how different brain regions activate in response to a painful stimulus. Moreover, there is considerable inter-subject variability in functional topology in pain response, causing brain locations to be misaligned across participants [35]. We anticipate these issues could lead to substantial issues in variable selection and model performance if outlying participants are not properly accounted for.
3 Assessing Inf luence on Variable Selection
We now characterize and define the influential observations on variable selection. It is critical to first recognize that observations influencing the choice of a submodel can only be legitimately understood within a stochastic framework due to the data-dependent nature of model selection. Moreover, due to distinctive model selection criteria, especially different penalty functions of regularized variable selection methods, contrasting sets of predictors could be chosen by different model selectors. Thus, it is reasonable to conjecture that the extent to which different model selectors are perturbed by influential points in terms of selection consistency properties varies. An implication is that influential points in our context are specific to model selectors, and the same detection procedure may well lead to the detection of distinctive sets of influential points.
In fact, the degree to which a model selector is prone to contamination can be termed as its functional sensitivity to influential observations. This concept can be understood as the percentage of data corruption permitted while fulfilling consistent selection properties. The significance of such sensitivity depends on the dichotomous intents of statistical operations, namely the diagnosis for contaminated data and consistent model selection. For diagnostic purposes, a model selector that is more sensitive to influential points may better facilitate their detection. On the other hand, the later objective of selection consistency underscores the resistance to contamination, which is closely linked to the notion of robustness of an estimator. Under this scenario, functional sensitivity philosophically aligns with the so-called breakdown point [16, 8] in robust statistics, which is the fraction of data allowed to be perturbed without invalidating the estimation consistency. From the above discussion, influential point set on variable selection is defined in a point-wise manner as:
Definition 1.
4 Methodology
@@
Let $\text{Y}\in\mathbb{R}$ be the response variable and ${x}_{j}\in\mathbb{R},\,j=1,...,p$, be the predictors. Let $\textbf{Y}=(\text{Y}_{1},...,\text{Y}_{n})^{\top}\in\mathbb{R}^{n}$ be the response vector and $\textbf{X}=(\textbf{X}_{1},...,\textbf{X}_{n})^{\top}\in\mathbb{R}^{n\times p}$, where $\textbf{X}_{i}=(\text{X}_{i1},...,\text{X}_{ip})^{\top}\in\mathbb{R}^{p}$, $i=1,...,n$, be the $n\times p$ design matrix. Then, the $i^{\text{th}}$ observation or point of dataset $\textbf{Z}\coloneqq(\textbf{Y},\textbf{X})$ is $\text{Z}_{i}\coloneqq(\text{Y}_{i},\textbf{X}_{i}^{\top})^{\top}\in\mathbb{R}^{p+1}$, $i=1,..,n$. Moreover, we abbreviate an arbitrary sequence of random variables or real values $a_{1},...,a_{n}$ by $a_{[n]}$.
@@
4.1 Single Inf luential Point Detection
The HIM and DF(LASSO) are revised in Sections 4.1.1 and 4.1.2, followed by proposing a generalized version in Section 4.1.3 .